Prev Next

SW&IT Meeting Minutes 2019-03-25

Attendees: Andrew, Igor, Matthieu, Petr, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every, Do not edit this table, use entries in smartsheets.

LR20190311 - Bec, Matthieu 81.0
LR20190311 - Colson, Andrew 80.0
LR20190311 - Edmeades, David 80.0
LR20190311 - Sola, Igor 32.0
LR20190311 - Kubanek, Petr 71.0
LR20190311 - Hooper, Stephen 80.0
LR20190311 - Zhang, Yang 80.0
LR20190311 - Zhang, Xianyu 80.0

News from management

Previous Action Items

Open topics

  • LBC (change from TFTP to NFS write of images)

Team reports

PK

next:

MB

  • OC, WPC, Management
  • LR automation in IMS, other minor improvements
  • iLocater, VoIP, OVMS+TT, L4AO, SSD19 meetings
  • meet the board, director's evaluation
  • timeseries db, continued experiment
  • extract pointing adjustements history from osa-nr into mongodb
next
  • timeseries db, continued experiment
  • generate group sub-systems responsibilty spreadsheet
  • follow-up VoIP (internally)

YZ

  • Some online safty training.
  • A new MODS IT#7610 assigned, will spend a bit of time studying it.
  • Recent LBC code change synced to github.
  • AO-Cravat:
    • Having some issues with github fork/merge, will fix it.
    • Still working on interface redesign.
    • Study Steve/Al's Python Strehl codes.
  • TCS/IIF/AOS:
    • Haven't received the name confirmation, but will go ahead and implement the "ApplyWaffleAO" in TCS/AOS.
    • Will go ahead and start instrument authorization code (consult Petr and other people).
    • Find the IIF document source and add the new command (low priority now).

AC

  • DIMM 3.4.3
    • added meanFlux value to DD
    • confirmed in DDBrowser and telemetry
    • discovered IIF.dimm telemetry data stream is obsolete; correct stream is DIMM.*
  • IT7612: looked into replacing misc weather info on FACSUM with radar image
  • Particle monitor
    • new error exceptions
      • telemetry exception: Unable to load leap table
      • initTelemetry() exception: This singleton has not been initialized
  • OAC testing today and tomorrow

SH

  • phones on the mountain -- meeting with UITS, MGIO
  • inventory, change from A-tags to S-tags.
  • worked on qemu-kvm integration on with VT-IO
  • archive and zabbix -- request of Elisa.
  • catalogs moved from the mountain to Tucson
  • Tucson VPN for various people
next:
  • Fedora 30 beta (maybe Atomic as well)
  • obs station gluster layout

IS

Last week
  • OT
    • worked on LMS tilted mask issue
    • Use the Telescope conversion scripts to transform and generate couple of programs for this ARGOS run
  • PIT
    • INAF pit work. Mostly done, need at least one more iteration with Michelle and work on PDF table
  • LUCI
    • LUCI issue tracks 7493
This week
  • PIT
    • Arizona spreadsheet
    • Maybe implementing anonymous proposals
    • Possibly some more fixes once the Italians come back with suggestions after their testing. Release is monday.
  • LUCI
    • Meet with Barry about issue 7473
    • Looking at implementing SRR mode for LUCI
    • Possibly installing latest GEIRS software??
  • OT
    • Investigate issue with failed script. AO and Guide stars coordinate mixing.
    • Do a script or application that wraps up the LMS tilted mask issue and allows easier querying.

XZ

  • OVMS+TT, WPC meetings
  • OVMS+TT test on the mountain
  • Review OVMS+TT related documents with Guido
  • Write OVMS+TT doccument and test plan
  • 4D SW follow up
next
  • OVMS+TT document
  • PBP action items
  • AO observation support
Topic revision: r13 - 25 Mar 2019, XianyuZhang
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback